OpenStudy (anonymous):

x^2+4x-5=0 is that 1, -4, 5, 4

OpenStudy (anonymous):

i thought it was 1

OpenStudy (cwrw238):

i assume you what are the zeroes of this function first factor this can u do that?

OpenStudy (anonymous):

is it like putting them together i have no clue??

OpenStudy (cwrw238):

factor x^2 + 4x - 5 = (x + 5)(x - 1) = 0 x-1 = 0 , x = 1 yes 1 is a root

OpenStudy (cwrw238):

x+ 5 = 0 x = -5 is the other root

OpenStudy (calculusfunctions):

I think the question is asking, which of the following are the roots of the given equation.

OpenStudy (cwrw238):

yes - so your answer is 1 teamjake

OpenStudy (anonymous):

ok i just wanted to check thanks

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

if there was an equation like x^2+x-6=0 i got -1 cause i did factor and it was -1 did i do it right?

OpenStudy (anonymous):

i thought i did

OpenStudy (cwrw238):

factors of this (x + 3)(x - 2)

OpenStudy (cwrw238):

the +x comes from 3x - 2x

OpenStudy (cwrw238):

zeros of this re 2, -3

OpenStudy (anonymous):

so -6

OpenStudy (cwrw238):